when Rs 105 is divided into two parts in the ratio of 2:5 , how much money will be in the first part?

Grade:9

2 Answers

Arun
3 years ago
Dear student
 
Let the part distributed = 2x and 5x
 
Hence
2x +5x = 105
 
7x = 105
 
x = 15
 
Hence first part = 2x = 2*15 = 30
MANTHAN RAO
3 years ago
ANSWER)30
EXPLAINATION)
LET ONE PART =X
THEREFORE 
FIRST PART =2X
SECOND PART=5X
SUM OF FIRST AND SECOND PART =105(AS 105 IS BEING DIVIDED)
2X+5X=105
7X=105
X=105/7
X=15
AS WE KNOW THAT FIRST PART IS 2X
THEREFORE FIRST PART =2*15
                                             =30
HENCE THE FIRST PART IS 30